OpenStudy (anonymous):

2. Use the trigonometric subtraction formula for sine to verify this identity: cos((π / 2) – x) = sin x

OpenStudy (anonymous):

@dpasingh

OpenStudy (blank ):

Is this a multiple choice question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@Blank 

OpenStudy (blank ):

cos(((π)/(2))-x)=sin(x) Use the difference formula for cosine to simplify the expression. The formula states that cos(A-B) = cosA cosB + sinA sinB. cos(((π)/(2)))*cos(x)+sin(((π)/(2)))*sin(x) Multiply all terms in the polynomial cos(((π)/(2)))cos(x)+sin(((π)/(2)))sin(x) Take the cosine of (π)/(2) to get 0. (0)(cos(x))+sin(((π)/(2)))sin(x) Multiply cos(x) by each term inside the parentheses. 0+sin(((π)/(2)))sin(x) Take the sine of (π)/(2) to get 1. 0+(1)(sin(x)) 0+1(sin(x)) 0+sin(x) sin(x)

OpenStudy (blank ):

Does that help?

OpenStudy (anonymous):

so what would I plug into those boxs?

OpenStudy (anonymous):

that helps a ton! But what answer would I put in each box?

OpenStudy (anonymous):

@Blank